OpenStudy (haleyelizabeth2017):

An object is dropped from a height of 1700 ft above the ground. The function h=-16t^2+1700 gives the object's height h in feet during free fall at t seconds. a. When will the object be 1000 ft above the ground? b. When will the object be 940 ft above the ground? c. What are a reasonable domain and range for the function h? a. I have about 6.61 seconds b. I have about 6.89 seconds c. I have no answer yet.....

TheSmartOne (thesmartone):

Your asnwer for a is correct.

TheSmartOne (thesmartone):

And so is the asnwer to B...

OpenStudy (haleyelizabeth2017):

okay.....so for c, do I plug in 0 for t and then solve for h? something like that?

TheSmartOne (thesmartone):

well the domain and range have to be positive...

TheSmartOne (thesmartone):

Becasue the height can't go negative and neither can the time...

OpenStudy (haleyelizabeth2017):

Okay. I get it now...so do I just solve: 16t^2=1700?

TheSmartOne (thesmartone):

Well the answer would just be domain is all postitive numbers and so is the range..

OpenStudy (aum):

Object is dropped from height h = 1700 ft. and eventually it will touch the ground (h = 0 feet). Range is: [0, 1700]

OpenStudy (confluxepic):

I got the same question.

TheSmartOne (thesmartone):

Oh yeah... oops

TheSmartOne (thesmartone):

Thanks for the catch..

OpenStudy (confluxepic):

I can still guide you.

OpenStudy (haleyelizabeth2017):

This is just confusing especially right now....:/ Are you in Algebra 2A confluxepic?

TheSmartOne (thesmartone):

And in that case domain is until y reaches 0

OpenStudy (confluxepic):

Algebra 2A Honors.

OpenStudy (haleyelizabeth2017):

me too!!!!

OpenStudy (haleyelizabeth2017):

so did you have the portfolio due today?

OpenStudy (confluxepic):

I am ahead by a week.

OpenStudy (haleyelizabeth2017):

oh :(

OpenStudy (confluxepic):

I did have a portfolio due today but I already did it before.

TheSmartOne (thesmartone):

So domain is from 0 to 10.31

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (haleyelizabeth2017):

so do I do what I put as the equation?

OpenStudy (aum):

Domain: \(0 \le t \le 10.31\) or [0, 10.31] Range: \(0 \le h \le 1700\) or [0, 1700]

OpenStudy (haleyelizabeth2017):

okay....thanks!

OpenStudy (aum):

yw.

OpenStudy (haleyelizabeth2017):

@confluxepic did you get #59? if so, how did you find how high in the air the ball goes?

OpenStudy (confluxepic):

What chapter and lesson? @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

I ended up figuring it out :/

OpenStudy (haleyelizabeth2017):

thank you though @confluxepic!